# Fast Flocking of Protesters on Street Networks

### v2

@@ Abstract
We propose a simple model of protesters scattered throughout a city who want to gather into large and mobile groups. This model relies on random walkers on a street network that follow tactics built from a set of basic rules. Our goal is to identify the most important rules for fast flocking of walkers. We explore a wide set of tactics and show the central importance of a specific rule based on alignment. Other rules alone perform poorly, but our experiments show that combining alignment with them enhances flocking.
Keywords : street networks, gathering, flocking, tactic design, protests
@@ 1 Introduction
Consider the following scenario. Protesters are scattered throughout a city and want to gather into groups large enough to perform significant actions. They face forces that may break up groups, block some places or streets and seize any communication devices protesters may be carrying. As a consequence, protesters only have access to local information on people and streets around them. Furthermore, formed protester groups must keep moving to avoid containment by adversary forces.
In this scenario, protesters need a distributed and as simple as possible protocol, that utilises local information exclusively and ensures the rapid formation of significantly large, mobile, and robust groups. We illustrate these objectives in Figure 1. Our goal in this paper is to identify the key building blocks for such protocols.
To do so, we model the city as a network of streets and intersections in Section 2. We then assume that protesters are biased random walkers on such a network. In Section 3, we present the set of basic rules that compose tactics walkers can use to move. We define, in Section 4, what flocking features groups must achieve and run extensive experiments to measure how effective each tactic is regarding those.
Figure 1: Walkers scattered on a street network (left) must rapidly gather (center) and subsequently flock (right).
@@ 2 Framework
We need a framework to simulate displacements of protesters in a city. We model cities as undirected graphs we call street networks. Protesters are then biased random walkers on this network. They can move from node to node with simple rules we introduce in next Section.
@@ 2.1 Street networks
In order to model real-world cities, we leverage OpenStreetMap [1] data and the OSMnx library [2]. For a given city, we use this library with its default settings to extract the graph $G=(V,E)$ defined as follows: the nodes in $V$ represent street intersections in this city and the links in $E\subseteq V\times V$ represent pieces of streets between them. We take the undirected graph $G$, meaning there is no distinction between $(u,v)$ and $(v,u)$ in $E$. In addition, we denote by $N(v)=\{u,(u,v)\in E\}$ the set of neighbors of any node $v$ in $V$.
We performed experiments on a wide ranges of large worldwide cities of diverse sizes and structures. This led to no significant difference on obtained results. We therefore use a typical instance, namely Paris, to present our work in this paper. This street network has 9 602 nodes, 14 974 links, leading to average degree 3.1. Its diameter is 83 hops and its average distance is 39.4 hops. The average street length is 99 meters, and the average distance is 5 552 meters.
The links of a street network generally represent street segments of very heterogeneous lengths [3]. Then, moves from a node to another one may have very different duration. In order to model this, we use a classical discretization procedure [4] that consists in splitting each link of the street network into pieces connected by evenly spaced nodes. In the obtained graph, each link represents a street slice of length close to a step $\delta$. Then, the walkers defined in previous section consistently make a move of length approximately $\delta$ at each time step.
In this paper, we use $\delta$ equal to 10 meters, leading to a network of $N=145000$ nodes and $M=300736$ links. It gives a sufficient precision in our context, and experiments with other reasonable values of $\delta$ displayed no significant difference.
@@ 2.2 Walkers
Given a network $G=(V,E)$, we consider a set $W$ of walkers numbered from $1$ to $|W|$. We denote the location of walker $i$ at time $t$ by $x_{i}(t)=v$, with $v\in V$. At each time step $t$, walker $i$ moves to node $x_{i}(t+1)\in N(v)$.
For any link $(u,v)$ in $E$, we also define the flux of walkers from $u$ to $v$ as $J_{u\rightarrow v}(t)=|\{i,x_{i}(t-1)=u,x_{i}(t)=v\}|$. We call group the set of walkers at a given node $v$ at a given time $t$: $g_{v}(t)=\{i,x_{i}(t)=v\}$. We denote by $n_{v}(t)=|g_{v}(t)|$ the number of walkers located at node $v$ at time $t$ and by $g(t)=|\{g_{v}(t),v\in V,g_{v}(t)\neq\emptyset\}|$ the number of non-empty groups at step $t$.
3 Walker model
@@ 3.1 Available information
We assume that walkers have very limited capabilities: they have no access to the actual location of other walkers, even on neighbor nodes, and they have no long-term memory and no communication protocol. Instead, we assume they only have access to estimates of aggregated observables, such as the number of walkers on neighbor nodes and the flow intensity on surrounding links.
More formally, walker $i$ may use the following information:
@@ 3.2 Criteria
  • Propulsion. The walker never goes back to its previous location $u$: the criterion has value 00 for $u$ and value $1$ for other neighbors of $v$.
  • Attraction. The walker preferably moves to nodes where there are already many walkers: the criterion is equal to the number of walkers $n_{w}(t)$.
  • Follow. The walker preferably follows the most popular moves of other walkers: the criterion has value $J_{v\rightarrow w}(t)$.
  • Alignment. The walker takes into account the net flux in both directions: the criterion is equal to $\Delta J_{v\rightarrow w}(t)=J_{v\rightarrow w}(t)-J_{w\rightarrow v}(t)$.
@@ 3.3 Walking rules
Parameter $\beta\geq 0$ is the intensity of choice: it quantifies the influence of the criterion on walker choices. If $\beta=0$, the criterion has no influence and walkers make purely random choices. If $\beta\rightarrow\infty$ walkers necessarily choose a neighbor among the ones that maximize the criterion.
3.4 Tactics
@@
where $\alpha_{C}$ is the coefficient of criterion $C$, with $\sum_{C\in\mathcal{C}}\alpha_{C}=1$. Therefore, a tactic is defined by a set of criteria and their coefficients. We call strict tactic one that has all its coefficients, except one, set to zero and then always follows the same criterion.
@@
As transition probabilities at time $t$ depend on previous non-deterministic moves, formal analysis of such processes is generally out of reach [5]. We will therefore explore possible tactics using simulations.
@@ 3.5 Baseline
More formally, for each node $v$ at time $t$, a unique random neighbor $u$ of $v$ is chosen and all walkers $i$, such that $x_{i}(t)=v$, move to $u=x_{i}(t+1)$. This is equivalent to purely random walks of groups until they meet another group.
@@
4 Experiments
We seek to design good tactics for our walkers. They must produce flocking faster than the baseline presented in Section 3.5. Moreover, we want to obtain short time convergence, as our walkers model the action of protesters that can not walk forever. In this Section we define metrics to evaluate how well walkers flock. We then run simulations for a reasonable duration and evaluate which tactics have performed the best in this time interval.
@@ 4.1 Flocking metrics
We characterize flocking as a gathering of walkers exploring the network. With notations of Section 2.2, this leads to the two following score definitions for a given run of a given tactic.
Definition 1 (cluster, gathering score $\rho(t)$).
A cluster is a maximal connected sub-graph with walkers on all its nodes. The gathering score $\rho(t)$ is the average number of walkers in clusters.
For example, when every agents are in the same cluster, $\rho(t)=|W|$. If instead all agents are in different ones, it equals $1$.
Definition 2 (mobility score $\mu(t)$).
@@
The mobility $\mu_{i}(t)$ of walker $i$ is the number of distinct nodes a walker has already visited at time $t$: $\mu_{i}(t)=|\{v,\exists t^{\prime}\leq t,x_{i}(t^{\prime})=v\}|$. The mobility score $\mu(t)$ is the average walker mobility.
@@
Notice that the mobility score is monotonically non-decreasing with time: $\mu(t+1)\geq\mu(t)$. In addition, if all walkers move to a node they already visited then $\mu(t+1)=\mu(t)$.
@@
We are interested in tactics with high gathering and mobility scores. Indeed, the high gathering score ensures that walkers form significant groups. In addition, the high mobility score implies that walkers continue to move. However, the fact that the gathering score remains high shows that walkers stay grouped. This means that the tactic successfully achieves flocking.
In order to gain more insight on group structure and dynamics, we introduce an additional metric.
Definition 3 (sprawling score $\sigma(t)$).
The sprawling score $\sigma(t)$ is the average number of nodes in clusters.
If the sprawling score is $1$, all groups are isolated from each other: whenever walkers are at a node, there is no walker on neighbor nodes. If instead the sprawling score is high, walkers form large clusters of neighbor groups. Its largest possible value is the total number of nodes in the network, meaning that there are walkers on each node.
We are interested in tactics with low sprawling score, meaning that they succeed in merging neighbor groups.
4.2 Extensive exploration of tactics
@@
In this Section, we follow an extensive method to explore the wide set of possible tactics on an entire city network:
@@
  • we consider the Paris street network discretized with parameter $\delta=10$ meters, leading to a network of $N=145000$ nodes and $M=300736$ links,
  • we consider $N$ walkers initially distributed uniformly at random on nodes,
@@
  • we perform $1000$ time steps, thus considering reasonably short walks of approximately $10$ kilometers,
  • we set $\beta$ large enough to ensure each walking rule strictly follows its criterion,
@@
  • we finally consider all tactics obtained as combinations of $\alpha_{C}$ parameter values from 00 to $1$ by steps of $0.1$ such that $\sum_{C\in\mathcal{C}}\alpha_{C}=1$.
@@
Figure 2: Gathering and mobility of all tactics. Each dot corresponds to the average last step value from ten runs of a tactic. The horizontal axis gives the gathering score, the vertical one gives the mobility score. From left to right and from top to bottom: tactics based mostly on the Random, Alignment, Attraction, Propulsion and Follow rule, respectively. On each of these plots, we indicate the strict tactic, that exclusively uses the corresponding rule. The bottom-right plot corresponds to tactics with no prevailing rule.
@@
With this setup, we obtain $1001$ different tactics. We run $10$ simulations of each tactic and we plot the average mobility and gathering scores in Figure 2. In these plots, each dot corresponds to a tactic defined by a set of $\alpha_{C}$ parameter values. We split these tactics into six plots: we display a set of tactics on the same plot if they all have $\alpha_{C}>0.5$ for the same criterion $C$, and we display on the last plot the set of all other tactics.
@@
We also display in each plot of Figure 2 a vertical and an horizontal line that indicate baseline results. Then, the tactics achieving the best flocking performances are the ones in the upper right corner: they obtain groups bigger and are more mobile that the baseline.
@@
We pay particular attention to strict tactics, which performances are spotted by red dots in Figure 2. We also highlight what we identify as the best tactic regarding our scores. It is the tactic with the highest gathering score among tactics that outperform our baseline. We display the evolution of mobility and gathering scores over time for these tactics in Figure 3.
@@
Figure 3: Plots showing the evolution of gathering, mobility and sprawling scores for the strict tactics and the best tactic. Gathering score is in log-log scale.
@@
Figure 2 clearly shows that Alignment-based tactics (top center plot) outperform others. All other sets of tactics perform poorly, except a few tactics for which no rule weights up more than 50% (bottom right plot). These tactics actually also use Alignment rule, to a lesser extent. This identifies the Alignment rule as a key building block for flocking tactics.

### v4

# Fast and Robust Flocking of Protesters in Street Networks

@@ Abstract
We present a simple model of protesters scattered throughout a city who want to gather into large and mobile groups. This model relies on random walkers on a street network that follow tactics built from a set of basic rules. Our goal is to identify the most important rules for fast and robust flocking of walkers. We explore a wide set of tactics and show the central importance of a specific rule based on alignment. Other rules alone perform poorly, but our experiments show that combining alignment with them enhances flocking, and that obtained groups are then remarkably robust.
Keywords: street networks, gathering, flocking, robustness, protests
@@ 1 Introduction
Consider the following scenario. Protesters are scattered throughout a city and want to gather into groups large enough to perform significant actions. They face forces that may break up groups, block some places or streets, seize communication devices carried by protesters, or even shut down or compromise communication systems. As a consequence, protesters only have access to local information on people and streets around them. Furthermore, formed protester groups must keep moving to avoid containment by adversary forces and perform actions.
In this scenario, protesters need a distributed and as simple as possible protocol, that uses local information only and ensures the fast formation of significantly large, mobile, and robust groups. We illustrate these objectives in Figure 1, and we call the expected collective behavior flocking. Our goal in this paper is to identify the key building blocks for flocking.
To do so, we first model the city as a network of streets and intersections, and we rely on deliberately strong simplifications to model protesters as biased random walkers on this network. We then define a set of basic rules that walkers may use as building blocks to compose tactics, as well as metrics that quantify key group features. We perform extensive experiments in order to explore how various tactics behave regarding these metrics. We also explore group robustness when adversary forces break them up while they follow an effective tactic. Finally, we discuss related work and highlight the originality of our contribution.
Figure 1: We expect walkers scattered on a street network (left) to quickly gather (center) and then flock (right).
@@ 2 Framework
This section presents our simulation framework. It models cities as undirected graphs that we call street networks. Then, protesters are biased random walkers on this network. They move from node to node following simple rules that we detail in next section. We keep our model as simple and minimal as possible in order to make it easy to explore the key principles behind flocking; our goal is not to model real-world protester behaviors, which is another important and challenging topic.
@@ 2.1 Street networks
In order to model real-world cities, we leverage OpenStreetMap data and the OSMnx library [4]. For a given city, we use this library with its default settings to extract the graph $G=(V,E)$ defined as follows: the nodes in $V$ represent street intersections in this city and the links in $E\subseteq V\times V$ represent the pieces of streets that exist between them. We focus on drivable streets because protesters generaly target them in order to maximize their impact on city routine. The graph $G$ is undirected, meaning there is no distinction between $(u,v)$ and $(v,u)$ in $E$. In addition, we denote by $N(v)=\{u,(u,v)\in E\}$ the set of neighbors of any node $v$ in $V$.
We performed experiments on a wide set of large worldwide cities of diverse sizes and structures. This led to no significant difference on obtained results. We therefore use a typical instance, namely Paris, to present our work in this paper. This street network has 9 602 nodes, 14 974 links, leading to average degree 3.1. Its diameter is 83 hops and its average distance is 39.4 hops  [^2] The distance between two nodes is the length of a shortest path between them. The average distance is the average of this length over all pairs pairs of nodes. The diameter is the maximal distance.. The average street length is 99 meters, and the average walking distance is 5 552 meters.
@@ 2.2 Walkers
Given a network $G=(V,E)$, we consider a set $W$ of walkers numbered from $1$ to $|W|$. We denote the location of walker $i$ at time $t$ by $x_{i}(t)=v$, with $v\in V$. At each time step $t$, walker $i$ moves to node $x_{i}(t+1)\in N(v)$. Rules we present in the following sections determine the choice of $x_{i}(t+1)$.
For any link $(u,v)$ in $E$, we also define the flux of walkers from $u$ to $v$ as $J_{u\rightarrow v}(t)=|\{i,x_{i}(t-1)=u,x_{i}(t)=v\}|$. We call group the set of walkers at a given node $v$ at a given time $t$: $g_{v}(t)=\{i,x_{i}(t)=v\}$. We denote by $n_{v}(t)=|g_{v}(t)|$ the number of walkers located at node $v$ at time $t$. We denote by $g(t)=|\{g_{v}(t),v\in V,g_{v}(t)\neq\emptyset\}|$ the number of non-empty groups at step $t$.
This definition allows groups of only one walker. We will see in Section 3.6 how to detect and avoid tactics that lead to such uninteresting groups.
2.3 Discretization
The links of a street network generally represent street segments of very heterogeneous lengths with heavy-tailed distributions [26]. Then, moving from an intersection to another one may have very different durations. In order to model this, we use a classical discretization procedure [28] that we describe below. It consists in splitting each link of the street network into pieces connected by evenly spaced nodes. We illustrate this in Figure 2.
More formally, given a link $(u,v)$ representing a piece of street of length $L$ in a street network and a discretization step $\delta$, we split each undirected link $(u,v)$ into $k$ new links, with $k=\lceil{\frac{L}{\delta}}\rceil$ as follows. We define a set of new nodes $V(u,v)=\{w_{0},w_{1},...,w_{k}\}$, the nodes from the original network being $w_{0}=u$ and $w_{k}=v$. We also define a set of new links $E(u,v)=\{(w_{i},w_{i+1}),i=0,1,...,k-1\}$. Then we define the set of all nodes $V=\bigcup_{u,v}V(u,v)$ and the set of links $E=\bigcup_{u,v}E(u,v)$. Finally, we model the street network by the graph $G=(V,E)$.
Figure 2: A piece of the discretized street network around Place de la Nation in Paris.
In the obtained graph, each link represents a street slice of length close to $\delta$. Then, the walkers defined in previous section consistently make a move of length approximately $\delta$ at each time step. In this paper, we use $\delta$ equal to 10 meters, leading to a network of n=130 276n=130\,276 nodes and m=300 736m=300\,736 links for Paris. It gives a sufficient precision in our context, and experiments with other reasonable values of $\delta$ displayed no significant difference.
3 Tactics
@@ 3.1 Available information
We assume that walkers have very limited capabilities. They have no long-term memory and no means of communication. They do not recognize other protesters and they do not have access to the location of other walkers, even on neighbor nodes. Instead, we assume they only have access to estimates of aggregated observables, such as the number of walkers on neighbor nodes and the flow intensity on surrounding links, as well as a very short-term memory.
More formally, walker $i$ may use the following information at time $t$:
@@ 3.2 Criteria
  • Propulsion. The walker never goes back to its previous location $u$, and otherwise makes random moves: the criterion has value 0 for $u$ and value $1$ for other neighbors of $v$.
  • Attraction. The walker preferably moves to nodes where there are already many walkers: the criterion is equal to the number of walkers $n_{w}(t)$ for each neighbor $w$ of $v$.
  • Follow. The walker preferably follows the most popular moves of other walkers: the criterion has value $J_{v\rightarrow w}(t)$ for each neighbor $w$ of $v$.
  • Alignment. The walker takes into account the net flux in both directions: the criterion is equal to $\Delta J_{v\rightarrow w}(t)=J_{v\rightarrow w}(t)-J_{w\rightarrow v}(t)$ for each neighbor $w$ of $v$.
Notice that the Alignment criterion may take negative values, when the flux towards $w$ is lower than the flux from $w$. This has a repulsion effect on walkers: they avoid rushing towards crowds that are already going to join them.
We chose these criteria because they are very simple and rely on local information only. They also model basic behaviors that may be observed in real-world protests, such as following others, avoiding going back, or joining already existing large groups. Last but not least, we also selected these criteria because they do not use any external information or communication, as wanted.
@@ 3.3 Walking rules
Parameter $\beta\geq 0$ is the intensity of choice: it quantifies the influence of the criterion on walker choices. If $\beta=0$, the criterion has no influence and walkers make purely random choices. If $\beta\rightarrow\infty$, walkers necessarily choose a neighbor among the ones that maximize the criterion.
This logit equation is widely used in the literature because it has many convenient features [6]. In particular, it ensures $\omega$ is a positive and monotonic function of the criterion: the greater $C_{u,v,w}(t)$, the greater the chances to move to node $w$ at time $t$. It is also compatible with criteria having both positive and negative values, like the flux-based one above.
We conducted experiments with other functions with these needed features; they led to no significant differences.
3.4 Rule aggregation
@@
where $\alpha_{C}$ is the coefficient of criterion $C$, with $\sum_{C\in\mathcal{C}}\alpha_{C}=1$. Therefore, a tactic is defined by a set of criteria and their coefficients.
We call strict tactic one that has all its coefficients, except one, set to zero and then always follows the same criterion.
@@
Notice that we may see a tactic as an inhomogeneous Markov chain, where transition probabilities are time-dependent. Still, at any given time, each walker makes a transition from its current state $(u,v)$ giving its previous location $u$ and its current location $v$, to its next state $(v,w)$ giving its current location $v$ and its next location $w$.
As transition probabilities at time $t$ depend on previous non-deterministic moves, formal analysis of such processes is generally out of reach [7]. We therefore explore possible tactics using simulations.
@@ 3.5 Baseline
More formally, for each node $v$ at time $t$, a unique random neighbor $u$ of $v$ is chosen and all walkers $i$ such that $x_{i}(t)=v$ move to $u=x_{i}(t+1)$. This is equivalent to purely random walks of groups until they meet another group.
@@ 3.6 Flocking metrics
In order to evaluate tactics, we need metrics describing how well walkers flock. We characterize flocking as a gathering of walkers that keep exploring the network.
This leads to the following score definitions for a given run of a given tactic. First, we define the coverage score, a classical concept to measure the efficiency of a walk for exploration [27].
Definition 3.1.
@@
The coverage $\kappa_{i}(t)$ of walker $i$ is the number of distinct nodes a walker has already visited at time $t$: $\kappa_{i}(t)=|\{v,\exists t^{\prime}\leq t,x_{i}(t^{\prime})=v\}|$. The coverage score $\kappa(t)$ is the average walker coverage: $\kappa(t)=\frac{1}{|W|}\sum_{i\in W}\kappa_{i}(t)$.
@@
Notice that the coverage score is monotonically non-decreasing with time and bounded by the number of nodes: $|V|\geq\kappa(t+1)\geq\kappa(t)$. In addition, if at time $t$ all walkers move to a node they already visited then $\kappa(t+1)=\kappa(t)$. In particular, walkers stuck in loops do not make the coverage score grow.
Second, we need a way to evaluate the ability of walkers to gather. The trivial approach consists in observing the average group size, a group being the set of walkers located at a same node. However, this makes no difference between two groups on neighbor nodes and two groups at very distant nodes. We therefore introduce the more advanced notion of clusters.
Definition 3.2.
We call cluster a maximal connected sub-graph with walkers on all its nodes. The node-size and the walker-size of a cluster is the total number of nodes involved in the cluster, and the total number of walkers on these nodes, respectively.
If all walkers are in the same cluster then the average walker-size of clusters is $\gamma(t)=|W|$. If instead all walkers are in different clusters, then the average walker-size of clusters is $\gamma(t)=1$. However, if there is a unique walker at each node of the considered graph, then there is a unique cluster and it contains all walkers. Then, the average walker-size of clusters is maximal, despite the fact that there is actually no significant group and no gathering. This is why we need to consider both the walker- and node-size of clusters, which lead to the following definition of gathering and sprawling scores.
Definition 3.3.
The gathering score $\gamma(t)$ is the average cluster walker-size at time $t$. The sprawling score $\sigma(t)$ is the average cluster node-size at time $t$.
@@
Since our main goal is to obtain large groups, we are primarly interested in tactics with high gathering score. Secondarily, we are interested in high coverage score and low sprawling score. Indeed, high gathering scores together with low sprawling scores ensure that walkers form significant clusters: many walkers are grouped on a small set of close nodes. A high coverage score implies in addition that walkers continue to move in the network. Therefore, the combination of such scores ensures that walkers move in groups, which means that the tactic successfully achieves flocking.
4 Exploration of tactics
We seek good tactics for our walkers. In particular, we seek tactics that produce flocking comparable to the baseline presented in previous section. Moreover, we want to obtain short time convergence, as our walkers model protesters that cannot walk forever. We therefore want tactics that produce flocking as fast as possible.
This section presents a wide set of simulations for realistic walk durations and explores which tactics perform best. We consistently provide interpretations of observed results and we deepen the study of the most effective tactics.
4.1 Behaviors on a single street
Our networks are composed of streets modeled as lines of nodes by the discretization presented above, see Figure 2. This plays an important role in walker dynamics, therefore we first detail this special case in this section.
Figure 3 displays space-time diagrams of three different strict tactics, i.e. the ones using only one rule, for $100$ walkers on a line of $100$ nodes. These space-time diagrams display the evolution of walker locations over time. The horizontal axis represents the nodes, numbered from $1$ to $100$. The vertical axis represents time steps, numbered from $1$ to $200$. Then, the color at coordinates $(x,y)$ represents the number of walkers at node $x$ after $y$ time steps, namely $n_{x}(y)$. The color is white if this number is 0, and it is the darkest if this number is the maximal value for the considered run. Walkers are initially uniformly distributed on all nodes.
Figure 3: Space-time diagrams of three strict tactics on a line using only (from left to right) the Random, Attraction, and Alignment rule. The fraction of walkers is the relative size of groups with respect to the maximum number of walkers in a group in a given run.
With the Random and with the Attraction strict tactics (Figure 3, left and middle), the largest obtained groups during the run both always contain less than 10 walkers (respectively 6 and 8 on the Figure) which never makes up to 10% of all walkers. However, the two tactics have distinct behaviors: while the Random one leads to walker constantly moving between groups, the Attraction one leads to stable groups. These groups fail to move along the line, though.
With Alignment (Figure 3, right), a more interesting pattern emerges: all walkers gather into a unique group that achieves flocking since it moves all along the line.
In conclusion, the Alignment strict tactic is sufficient to obtain flocking on a street. However, the groups we obtain eventually reach the street extremities, which raises the question of their behavior at street intersections.
4.2 Extensive exploration
@@
In this section, we use an extensive method to explore the wide set of possible tactics on an entire city network:
@@
  • we consider the Paris street network discretized with parameter $\delta=10$ meters, leading to a network of $n=130276$ nodes and $m=300736$ links,
  • we consider a set $W$ of $|W|=n$ walkers initially distributed uniformly at random in the network,
@@
  • we perform $1000$ time steps, thus considering reasonably short walks of approximately $10$ kilometers at most,
@@
  • we consider all tactics obtained as combinations of $\alpha_{C}$ parameter values from 0 to $1$ by steps of $0.1$ such that $\sum_{C\in\mathcal{C}}\alpha_{C}=1$,
  • finally, we set $\beta$ large enough to ensure that each walking rule strictly follows its criterion. Indeed, in the following experiments, lowering this value only worsened gathering scores, as it is equivalent to adding noise to walkers choices.
@@
Figure 4: Gathering and coverage scores of all tactics. Each dot corresponds to the average last step value from ten runs of a tactic. The horizontal axis gives the gathering score, the vertical one gives the coverage score. From left to right and from top to bottom: tactics based mostly on the Random, Alignment, Attraction, Propulsion and Follow rule, respectively. On each of these plots, the red dot is for the strict tactic, that exclusively uses the corresponding rule. The bottom-right plot corresponds to tactics with no prevailing rule.
@@
With this setup, we obtain $1001$ different tactics. We run $10$ simulations of each tactic and plot the average coverage and gathering scores in Figure 4. In these plots, each dot corresponds to a tactic, defined by a set of $\alpha_{C}$ parameter values. We split these tactics into six plots: we display a set of tactics on the same plot if they all have $\alpha_{C}>0.5$ for the same criterion $C$, and we display on the last plot the set of all other tactics.
@@
We also display in each plot of Figure 4 vertical and a horizontal lines that indicate baseline results. Then, the tactics achieving the best flocking performances are the ones in the upper right corner: they obtain larger and more mobile groups than the baseline.
@@
We pay particular attention to strict tactics, which performances are spotted by red dots in Figure 4. We also highlight what we identify as the best tactic regarding our scores. It is the tactic with the highest gathering score among tactics that outperform our baseline. For more insight, we display the evolution of coverage, gathering and sprawling scores over time for these tactics in Figure 5.
First notice that all strict tactics have very poor scores, except the Alignment one. This is a consequence of our previous section results, where it was the only one that generates large and mobile groups. Other strict tactics failed to achieve those patterns in a street, and therefore fail on an entire city.
Attraction, Follow and Random strict tactic have almost no dynamic, as walkers do not explore much the network and do not form large groups. Propulsion however displays an excellent coverage score with an almost linear growth on Figure 5, as walkers never go back to their previous location.
Note that tactics in the scatter plot (Figure 4) are not uniformly distributed. Indeed, no tactic has gathering and coverage scores such that its corresponding dot would be in the bottom right corner of the plot. This shows that a high coverage score is a necessary condition to achieve a high gathering score, but it is not sufficient.
@@
Figure 5: Plots showing the evolution of gathering, coverage and sprawling scores for the strict tactics and the best tactic. Gathering score is in log-log scale for readability.
@@
Figure 4 clearly shows that Alignment-based tactics (top right plot) outperform others. All other sets of tactics perform poorly, except a few tactics for which no rule weights up more than 50% (bottom right plot). These tactics actually also use Alignment rule, to a lesser extent. This identifies the Alignment rule as a key building block for flocking tactics.